The Greenback Stores cost structure is dominated by variable costs with a contribution margin ratio of 0.25 and fixed costsSuppose that both companies experience a 15 percent increase in sales volume. By how much would each companys profits increa
Answer #1

Question 1

Greenback Store One Mart
Particulars Amount Percentage Amount Percentage
Sales 800,000 100% 800,000 100%
Variable Costs 600,000 75% 200,000 25%
Contribution Margin 200,000 25% 600,000 75%
Fixed Costs 40,000 5% 440,000 55%
Operating Profit 160,000 20% 160,000 20%

Notes

Greenback Store

Variable Costs % = 100% - Contribution Margin Percentage

Variable Costs % = 100% - 25% = 75%

Variable Costs = 75% of $ 800,000 = $ 600,000

Operating Profit = Sales - Variable Cost - Fixed Costs

For One Mart

Variable Costs % = 100% - Contribution Margin Percentage

Variable Costs Percentage = 100% - 75% = 25%

Variable Costs = 25% of $ 800,000 = $ 200,000

Operating Profit = Sales - Variable Costs - Fixed Costs

Additional Note

Fixed Costs Percentage = Fixed Costs / Sales * 100

Operating Profit Percentage = Operating Profit / Sales * 100

Question 2

For Green Mart

Degree of Operating Leverage = Contribution Margin / Operating Income

Contribution Margin = $ 200,000

Operating Income = $ 160,000

Degree of Operating Leverage = 200,000 / 160,000 = 1.25 Times

Now Calculation of Increase in Profit with 15% Increase in Sales

Degree of Operating Leverage = % Change in Operating Income / % Change in Sales

1.25 = % Change in Operating Income / 15%

% Change in Operating Income = 1.25 * 15% = 18.75%

Increase in Income = Income before Change in Sales * % Change in Operating Income

Increase in Income = 160,000 * 18.75%

Increase in Income = $ 30,000

For One Mart

Degree of Operating Leverage = Contribution Margin / Operating Income

Contribution Margin = $ 600,000

Operating Income = $ 160,000

Degree of Operating Leverage = 600,000 / 160,000 = 3.75 Times

Now Calculation of Increase in Profit with 15% Increase in Sales

Degree of Operating Leverage = % Change in Operating Income / % Change in Sales

3.75 = % Change in Operating Income / 15%

% Change in Operating Income = 3.75 * 15% = 56.25%

Increase in Income = Income before Change in Sales * % Change in Operating Income

Increase in Income = 160,000 * 56.25%

Increase in Income = $ 90,000